A contact lens is made of plastic with an index of refraction of 1.40. The lens has an outer radius of curvature of +1.98 cm and an inner radius of curvature of +2.52 cm. What is the focal length of the lens?
_____ cm
Answer #1

According to Lens Maker's formula,

1/f = (n-1) (1/R1 - 1/R2)

=> 1/f = (1.40-1) ( 1/1.98 - 1/2.52)

=> 1/f = 0.0433

=> f = 23.09 cm